VII. The photoelectric threshold wavelength of a tungsten surface is 272 nm. Calculate the maximum
kinetic energy of the electrons ejected from this tungsten surface by ultraviolet radiation of frequency
1.45 x 1015 Hz. Express the answer in electron volts.

Notes:
Remember that E=hf=hc/λ
Threshold wavelength means the minimum amount of wavelength to remove electrons from a surface
so in this problem, E = = hf ϕ threshold=hc/ λthreshold
where λthreshold = threshold wavelength
= work function = the minimum energy needed to remove an electron from the surface ϕ
h = Planck’s constant = 6.626 x 10 -34 J.s = 4.136 x 10 -15 eV.s
c = 3x108 m/s
f = frequency.
1
Expert's answer
2020-06-08T10:29:45-0400

"\frac{hc}{\lambda}=A+KE"


"A=\frac{hc}{\lambda_0}=\frac{6.62\cdot 10^{-34}\cdot 3\cdot10^8}{272\cdot10^{-9}}=7.3\cdot10^{-19}J"


"KE=h\nu-7.3\cdot10^{-19}=6.62\cdot 10^{-34}\cdot 1.45\cdot10^{15}-7.3\cdot10^{-19}="


"=2.3\cdot10^{-19}J=1.44eV"